Thomson Media is considering some new equipment whose data are shown below. The equipment has a 3-year tax life and would be fully depreciated by the straight-line method over 3 years, but it would have a positive pre-tax salvage value at the end of Year 3, when the project would be closed down. Also, additional net operating working capital would be required, but it would be recovered at the end of the project's life. Revenues and other operating costs are expected to be constant over the project's 3-year life. What is the project's NPV? Do not round the intermediate calculations and round the final answer to the nearest whole number.

WACC 10.0%

Net investment in fixed assets (depreciable basis) $70,000

Required net operating working capital $10,000

Straight-line depreciation rate 33.333%

Annual sales revenues $57,000

Annual operating costs (excl. depreciation) $30,000

Expected pre-tax salvage value $5,000

Tax rate 35.0%
Answer #1
Calculation of Project's NPV
Particulars 0 1 2 3
Initial Investment:
Investment in Fixed Assets -70000
Net Working Capital investment -10000
Net Initial Investment (A) -80000
Operating Cash Flows:
Annual Sales Revenue (B) 57000 57000 57000
Less: Annual Operating Costs (C ) 30000 30000 30000
Less: Depreciation (D)
($70,000 * 33.333%)
23333.1 23333.1 23333.1
Profit Before Tax (E = B-C-D) 3666.9 3666.9 3666.9
Less: Tax @35% (F = E*35%) 1283.415 1283.415 1283.415
Profit After Tax (G = E-F) 2383.485 2383.485 2383.485
Add back Depreciation (H = D) 23333.1 23333.1 23333.1
Net Operating Cash Flows (I = G+H) 25716.59 25716.59 25716.59
Terminal Value:
Pre tax Salvage value (J) 5000
Less: Tax @35% (K) 1750
After Tax Salvage Value (L = J-K) 3250
Net working Capital Recaptured (M) 10000
Net Terminal value (N = L+M) 13250
Total Cash Flows (O = A+I+N) -80000 25716.59 25716.59 38966.59
Discount Factor @10% (P)
1/(1+10%)^n
n=0,1,2,3
1 0.909091 0.826446 0.751315
Discounted Cash Flows (Q = O*P) -80000 23378.71 21253.38 29276.17
NPV of the Project -6091.74
Therefore, Project's NPV is ($6,091.74)